# Membrane phononic integrated circuits

### v1

@@ VII Coupled mode theory
CMT is generally applicable to situations involving waves at interfaces. It derives only from conservation of energy and the approximation of ‘small coupling’, which assumes that to first order the presence of the coupling does not modify the eigenmodes of the system [56]. A key property of CMT is that the change in amplitude of one of the coupled elements is proportional to the amplitude of motion of the other coupled element [23]. The constant of proportionality is called the coupling rate. For example if $a$ is the amplitude of energy in the eigenmode of interest, which is coupled to $N$ other modes, we can write:
@@
Here $\gamma_{i}$ and $a_{i}$ are the coupling rate and amplitude of the $i^{\mathrm{th}}$ external eigenmode. We usually normalise the amplitudes so that $|a|^{2}$ equals the energy in the mode of interest and $|a_{i}|^{2}$ equals the energy in the $i^{\mathrm{th}}$ eigenmode.
@@
Figure 22: Origin of work and power transfer in phononic coupled mode theory. (a) We consider an eigenmode (blue, above) with time-varying displacement $u_{L}$, localised to a region of increased density $\rho_{L}>\rho_{0}$ (green, below). Without external coupling, a volume element in the region of lower density (where the field is evanescent) experiences a force per unit volume of $\rho_{0}\ddot{u}_{L}$. The net work associated with this force is zero, with the force simply shifting energy between potential and kinetic forms. (b) We perturbatively introduce a second eigenmode (orange), with its own time-varying displacement $u_{R}$, localised to a region of increased density $\rho_{R}$. The force on a volume element in this region, due to the left (blue) eigenmode, is $(\rho_{R}-\rho_{0})\ddot{u}_{L}$. Consequently, the net flow of power from the left mode to the right mode (red arrow) is $(\rho_{R}-\rho_{0})\ddot{u}_{L}\dot{u}_{R}$. Depending on the relative phase of the oscillations, this power may be positive or negative.
The origin of coupling in phononic CMT is illustrated in Figure 22. In Fig. 22(a) we consider an eigenmode (blue shading) with a time-varying displacement $u_{L}$, localised to a region of increased density $\rho_{L}>\rho_{0}$, where $\rho_{0}$ is the background density. In the surrounding material with density $\rho_{0}$ the acoustic field from the mode is evanescent. However, the nonzero amplitude still exerts a force per unit volume of $\rho_{0}\ddot{u}_{L}$. In the absence of coupling, the net work associated with this force is zero: it simply converts energy between kinetic and potential forms as the membrane vibrates. In Fig. 22(b) we introduce a second eigenmode (orange shading) with time-varying displacement $u_{R}$, localised to a region of increased density $\rho_{R}>\rho_{0}$. Under the perturbative approach of the coupled-modes formalism, the introduction of the $u_{R}$ mode does not affect the $u_{L}$ mode. Therefore, we can say that the force on a volume element in the $u_{R}$ mode, due to the vibration of the $u_{L}$ mode, is $(\rho_{R}-\rho_{0})\ddot{u}_{L}$. This has an associated power per unit volume from the $u_{L}$ mode to the $u_{R}$ mode, $P_{LR}$ (red arrow), of:
@@ VII.1 Analytic expression for coupling rates
We will use Eq. (76) to derive an analytic expression for the coupling rates between evanescently coupled phononic resonators and waveguides. We restrict attention to evanescent coupling because CMT, as a perturbative treatment, is only valid when the coupling is relatively weak (specifically, when it does not change the shape of the modes).
@@
Our system of interest is two parallel, evanescently coupled waveguides, as illustrated in Fig. 23(a). The two waveguides are defined by local increases in volumetric density (green shading): $\rho_{L}$ on the left and $\rho_{R}$ on the right, both being greater than the background density $\rho_{0}$. Superimposed on the waveguide densities we have plotted the fundamental eigenmodes of each waveguide (blue and orange curves), calculated from Eq. (74) (see caption for parameter values). The eigenmodes both have evanescent components which will produce the coupling. The next subfigures provides examples for how this density modification could be fabricated: by reducing the density with a patterned mesh outside the waveguiding region (Fig. 23(b)); by increasing density with a ridge acoustic waveguide set within the mesh (Fig. 23(c)); and by increasing the density by depositing material such as metal (Fig. 23(d)).
@@
where $\psi_{L}(x)$ is the dimensionless function in the range [1,1][-1,1] that describes the eigenmode from Eq. (70).
In the same way that we derived Eq. (76), we can say that a volume element in the right waveguide experiences a net power $P_{LR}$ due to the motion in the left waveguide, equal to:
$P_{LR}=(\rho_{R}-\rho_{0})\ddot{u_{L}}\dot{u}_{R}.$ (80)
Evaluating the time derivatives will pull out a factor of $\Omega^{3}\cos(\Omega t)\sin(\Omega t+\Delta\varphi)$, which on average equals $\frac{1}{2}\Omega^{3}\sin(\Delta\varphi)$. Therefore the time averaged power per unit volume flowing from the left waveguide to the right is:
@@
The $\sin(\Delta\varphi)$ factor is significant as it determines the direction of power flow. The greatest power flow occurs when $|\Delta\varphi|=\pi/2$ and no power flow occurs when $|\Delta\varphi|=0\;\text{or}\;\pi$. This resembles the spectral response of a driven lumped-element oscillator (see Fig. 35 from Section IV.2); in that case the greatest response occurs when the drive is $\pi/2$ ahead in phase with the resonator. Here the left waveguide does work on the right waveguide when $\Delta\varphi$ is in the first or second quadratures, corresponding to the waves in the left waveguide appearing ‘ahead’ of the waves in the right waveguide. Vice versa holds when power flows in the opposite direction. We see this borne out later in Figure 24(a).
@@
The integral here is over the right waveguide (i.e. over everywhere where, if the right waveguide was ‘thought away’, the density would be $\rho_{0}$).
Now we will pause at Eq. (VII.1) and derive the same quantity using CMT. The coupled modes description of the same two waveguides is a pair of coupled differential equations:
@@
Substituting Eq.s (83) into Eq. (84) and simplifying will yield that $\gamma_{RL}=-\gamma_{LR}^{*}$, which also means that $|\gamma_{LR}|^{2}=|\gamma_{RL}|^{2}=|\gamma|^{2}$.
From Eq.s (83) we can derive the rate of change of $|a_{R}(y)|^{2}$ (we omit arguments of $y$ for clarity):
$\frac{\mathrm{d}|a_{R}^{2}|}{\mathrm{d}y}=a_{R}^{*}\frac{\mathrm{d}a_{R}}{\mathrm{d}y}+a_{R}\frac{\mathrm{d}a_{R}^{*}}{\mathrm{d}y}=\gamma_{LR}a_{L}a_{R}^{*}+\gamma_{LR}^{*}a_{L}^{*}a_{R}.$ (85)
Here we have used the relationship that $\gamma_{LR}=\gamma_{RL}^{*}$.
@@
This can be heuristically justified by taking the real part of Eq. (85), which produces a factor of $2$, or by comparing to very similar calculations in the literature [119, 23]. In the next section we will also quantitatively justify this expression with a numerical simulation. The symmetric equation for coupling into the left waveguide is:
@@
A key insight of Eq.s (86) and (87) is that the coupling rate is proportional to the overlap integral of the normalised modeshapes. This immediately provides an intuitive explanation for the coupling of modes. We can see that the modeshapes must be spatially proximate to achieve coupling. Additionally, because the overlap integral is a signed integral, modes with differing even and odd symmetry will couple with each other minimally. We will use this second principle later (see Section XI.4) to design a mode demultiplexing power splitter.
@@ VII.2 Numerical test
Figure 24: Finite difference time domain (FDTD) simulation of evanescent coupling between two parallel waveguides. Waves are launched at the left side of the simulation at frequency $\Omega/2\pi=12\,\mathrm{MHz}$. The heatmap plots the displacement in arbitrary units, after sufficient time such that waves have propagated across the length of the waveguide. (a) Coupling between two identical waveguides. The waveguides have width $30\,\mathrm{\upmu m}$ and separation $20\,\mathrm{\upmu m}$. Insets (i) and (ii) highligt the phase lag predicted by Eq. (81), which changes sign with the direction of power transfer. (b) Coupling in a nearly identical simulation where the width of the non-drive waveguide is increased to $40\,\mathrm{\upmu m}$. Parameters: the stress $\sigma$ = 1 GPa, the density in both waveguides is $\rho=3100\,\mathrm{kg\cdot m^{-3}}$, and the density outside the waveguides is $\rho/4$.
We begin by returning to Eq.s (83). If we assume the slowly-varying amplitudes change with $y$ as $e^{iky}$ for some $k$, then the equations can be rearranged as:
@@
Given initial conditions $a_{L}(0)$ and $a_{R}(0)$ and our assumptions of codirectional and positive group velocities, Eq.s (83) can be solved for the two amplitudes [56] to give:
@@
These equations are mathematically equivalent to those that govern optical couplers and switches [56, 121]. The power in each waveguide oscillates with a wavelength of $\lambda_{\mathrm{trans}}=2\pi/k_{c}$, which in terms of coupling rates is equal to:
@@
If we assume that initially all the power is in one waveguide, say $a_{R}(0)=0$, then the maximum amount of power is in waveguide $R$ when $\cos(k_{c}y)=0$ and $\sin(k_{c}y)=\pm 1$. This occurs for example at $y=\pi/2k_{c}$, implying the maximum amount of power that can be transferred is:
@@
The results of the simulation and our coupled modes expression are summarised in Table 1. The relative error between the analytic theory and simulation is generally $\lesssim 10\%$, and we can observe the qualitative behaviour predicted by the theory. Specifically, we see oscillation of power between the two waveguides. The two insets highlight the phase lag between the wavefronts predicted by Eq. (81). We can see for example in inset (i) that the bottom waveguide appears ‘ahead’ of the top waveguide at the time where there is a net power flow from it to the top waveguide.
@@ VIII Longitudinal evanescent coupling
A significant advantage of tunnel couplers is that they are directly compatible with highly-acoustically confining, highly-impedance mismatched suspended membranes. In contrast, coupling schemes like the parallel waveguides example in Section VII.1 rely on the lateral evanescent field and therefore must have a comparatively low impedance contrast. To achieve such a low impedance mismatch with suspended membranes would require a more difficult fabrication process and quite possibly perform worse—for example, waveguides could be defined by depositing metal but this would introduce surface losses [24]. In other words, evanescent tunnel couplers are the natural solution for creating coupling in suspended membrane devices. They can also be used to create resonators with high (and tailorable) quality factors, as we will now discuss.
@@ IX Input-output formalism
Here $\gamma_{1}$ and $\gamma_{2}$ (with dimension $T^{-1}$) are the coupling rates to waveguides $1$ and $2$ respectively, and $\gamma_{0}$ is the intrinsic or internal loss rate describing coupling to the environment. The total loss rate $\gamma$ is equal to the sum of the couplings to the waveguides and the environment: $\gamma=\gamma_{0}+\gamma_{1}+\gamma_{2}$. The situation can be generalised to resonators coupled to $N$ waveguides by considering additional coupling rates $\gamma_{3},\,\gamma_{4},\,\ldots$ and so on. If the resonator is not being driven we can easily derive from Eq. (96) that the rate of energy loss is $\frac{\mathrm{d}|a|^{2}}{\mathrm{d}t}=-\gamma|a|^{2}$, consistent with our previous experience that $\gamma/2$ is the rate of amplitude decay.
@@
As illustrated in Fig. 28, the acoustic field in each waveguide is a superposition of waves travelling towards and away from the resonator. Because we assume the waveguides are reflectionless, any waves travelling towards the resonator ($a_{\mathrm{{in,1}}},\,a_{\mathrm{{in,2}}}$) must originate from external sources; in contrast waves travelling away can either come from the resonator or be reflected inputs. It can be shown by applying time reversal symmetry that the loss rate for each waveguide also defines its reflection coefficient [131], such that the outgoing amplitude in each waveguide is simply:

### v2

@@ VII Coupled mode theory
CMT is generally applicable to situations involving waves at interfaces. It derives only from conservation of energy and the approximation of ‘small coupling’, which assumes that to first order the presence of the coupling (i.e. perturbative) element does not modify the eigenmodes of the system [63]. A key property of CMT is that the change in amplitude of one of the coupled elements is proportional to the amplitude of motion of the other coupled element [23]. The constant of proportionality is called the coupling rate. For example if $a$ is the amplitude in the eigenmode of interest, which is coupled to $N$ other modes, we can write:
@@
Here $\gamma_{i}$ and $a_{i}$ are respectively the coupling rate to and amplitude of the $i^{\mathrm{th}}$ external eigenmode. We usually normalise the amplitudes so that $|a|^{2}$ equals the energy in the mode of interest and $|a_{i}|^{2}$ equals the energy in the $i^{\mathrm{th}}$ eigenmode.
@@
Figure 22: Origin of work and power transfer in phononic coupled mode theory. (a) We consider an eigenmode (blue, above left) with time-varying displacement $u_{L}$, localised to a region of increased density $\rho_{L}>\rho_{0}$ (green, below left). Without external coupling, a volume element in the region of lower density (where the field is evanescent) experiences a force per unit volume of $\rho_{0}\ddot{u}_{L}$. The net work associated with this force is zero, with the force simply shifting energy between potential and kinetic forms. (b) We perturbatively introduce a second eigenmode (orange, right), with its own time-varying displacement $u_{R}$, localised to a region of increased density $\rho_{R}$. The force on a volume element in this region, due to the left (blue) eigenmode, is $\rho_{R}\ddot{u}_{L}$. Consequently, the net flow of power from the left mode to the right mode (red arrow) associated with this volume element is $(\rho_{R}-\rho_{0})\ddot{u}_{L}\dot{u}_{R}$. Depending on the relative phase of the oscillations, this power flow may be positive or negative.
The origin of coupling in phononic CMT is illustrated in Figure 22. In Fig. 22(a) we consider an eigenmode (blue shading) with a time-varying displacement $u_{L}$, localised to a region of increased density $\rho_{L}>\rho_{0}$, where $\rho_{0}$ is the background density. In the surrounding material with density $\rho_{0}$ the acoustic field from the mode is evanescent. However, the nonzero amplitude still exerts a force per unit volume of $\rho_{0}\ddot{u}_{L}$. In the absence of coupling, the net work associated with this force is zero: it simply converts energy between kinetic and potential forms as the membrane vibrates. In Fig. 22(b) we introduce a second eigenmode (orange shading) with time-varying displacement $u_{R}$, localised to a region of increased density $\rho_{R}>\rho_{0}$. Under the perturbative approach of the coupled-modes formalism, the introduction of the $u_{R}$ mode does not affect the $u_{L}$ mode. Therefore, we can say that the force on a volume element in the $u_{R}$ mode, due to the vibration of the $u_{L}$ mode, is $\rho_{R}\ddot{u}_{L}$. The associated instantaneous power per unit volume (proportional to the force times velocity) from the $u_{L}$ mode to the $u_{R}$ mode, $P_{LR}$ (red arrow) is:
@@ VII.1 Analytic expression for coupling rates
We will use Eq. (77) to derive an analytic expression for the coupling rates between evanescently coupled phononic resonators and waveguides. We restrict our attention to evanescent coupling because CMT, as a perturbative treatment, is only valid when the coupling is relatively weak (specifically, when it does not change the shape of the modes).
@@
Our example system of interest is two parallel, evanescently coupled waveguides, as illustrated in Fig. 23(a). The two waveguides are defined by local increases in volumetric density (green shading): $\rho_{L}$ on the left and $\rho_{R}$ on the right, both being greater than the background density $\rho_{0}$. Superimposed on the waveguide densities we have plotted the fundamental eigenmodes of each waveguide (blue and orange curves), calculated from Eq. (75) (see caption for parameter values). The eigenmodes both have evanescent components which will produce the coupling. The next subfigures provides examples for how this density modification could be fabricated: by reducing the density with a patterned mesh outside the waveguiding region (Fig. 23(b)); by increasing density with a ridge acoustic waveguide set within the mesh (Fig. 23(c)); and by increasing the density by depositing material such as metal (Fig. 23(d)).
@@
where $\psi_{L}(x)$ is the dimensionless function with $\max(|\psi_{L}(x)|)=1$ that describes the eigenmode from Eq. (71).
A volume element in the right waveguide experiences a net power $P_{LR}$ due to the motion in the left waveguide according to Eq. (77). Substituting $u_{L}$ and $u_{R}$ with Eqs. (78) and (79), then evaluating the time derivatives will pull out a factor of $\Omega^{3}\cos(\Omega t)\sin(\Omega t+\Delta\varphi)$, which on average equals $\frac{1}{2}\Omega^{3}\sin(\Delta\varphi)$. Therefore the time averaged power per unit volume flowing from the left waveguide to the right is:
@@
The $\sin(\Delta\varphi)$ factor is significant as it determines the direction of power flow. The greatest power flow occurs when $|\Delta\varphi|=\pi/2$ and no power flow occurs when $|\Delta\varphi|=0\;\text{or}\;\pi$. This resembles the spectral response of a driven lumped-element oscillator (see Fig. 9 from Section IV.2); in that case the greatest response occurs when the drive is $\pi/2$ ahead of the resonator in phase. Here the left waveguide does work on the right waveguide when $\Delta\varphi$ is in the first or second quadratures, corresponding to the waves in the left waveguide appearing ‘ahead’ of the waves in the right waveguide. Vice versa holds when power flows in the opposite direction. We see this borne out later in finite difference time domain simulations shown in Figure 24(a).
@@
The integral here is over the right waveguide (i.e. where in Fig. 23(a) the density has increased from $\rho_{0}$ to $\rho_{R}$).
Now we will pause at Eq. (VII.1) and derive the same quantity using CMT. The coupled modes description of the same two waveguides is a pair of coupled differential equations [23]:
@@
Substituting Eq. (83) into Eq. (84) and simplifying will yield that $\gamma_{RL}=-\gamma_{LR}^{*}$, which also means that $|\gamma_{LR}|^{2}=|\gamma_{RL}|^{2}=|\gamma|^{2}$.
From Eq. (83) we can derive the rate of change of $|a_{R}(y)|^{2}$ (we omit arguments of $y$ for clarity):
$\frac{\mathrm{d}|a_{R}|^{2}}{\mathrm{d}y}=a_{R}^{*}\frac{\mathrm{d}a_{R}}{\mathrm{d}y}+a_{R}\frac{\mathrm{d}a_{R}^{*}}{\mathrm{d}y}=\gamma_{LR}a_{L}a_{R}^{*}+\gamma_{LR}^{*}a_{L}^{*}a_{R}.$ (85)
@@
This can be heuristically justified by taking the real part of Eq. (85), or by comparing to similar calculations in the literature derived in the case of evanescent coupling of electromagnetic waves [128, 23]. In the next section we will also quantitatively justify this expression with a numerical simulation. The symmetric equation for coupling into the left waveguide is:
@@
A key insight of Eqs. (86) and (87) is that the coupling rate is proportional to the overlap integral of the normalised modeshapes. This immediately provides an intuitive explanation for the coupling of modes. We can see that the modeshapes must be spatially proximate to achieve coupling. Additionally, because the overlap integral is a signed integral, modes with differing even and odd symmetry will couple with each other minimally. We will use this second principle later (see Section XI.4) to design a mode demultiplexing power splitter.
@@
We note that this CMT approach is quite general and can be applied as well to model the coupling of Rayleigh-like or Love-like acoustic modes in on-chip high acoustic index contrast slab waveguides [3, 50, 49, 48, 5]. The specific case of coupling of flexural modes of acoustic membranes detailed here provides simple expressions (Eqs. (86) and (87)) reminiscent of the results obtained in the optical domain [23, 128, 130], as the displacement is purely one dimensional (out of plane) and does not involve any material anisotropy [5]
@@ VII.2 Numerical test
Figure 24: Finite difference time domain (FDTD) simulation of evanescent coupling between two parallel waveguides. Waves are launched at the left side of the simulation at frequency $\Omega/2\pi=12\,\mathrm{MHz}$. The heatmap plots the displacement in arbitrary units, after sufficient time such that waves have propagated across the length of the waveguide. (a) Coupling between two identical waveguides. The waveguides have width $30\,\mathrm{\upmu m}$ and separation $20\,\mathrm{\upmu m}$. Insets (i) and (ii) highlight the phase lag predicted by Eq. (81), which changes sign with the direction of power transfer. (b) Coupling in a nearly identical simulation where the width of the non-drive waveguide is increased to $40\,\mathrm{\upmu m}$. Parameters: stress $\sigma$ = 1 GPa; density in both waveguides is $\rho=3100\,\mathrm{kg\cdot m^{-3}}$; and density outside the waveguides is $\rho/4$.
We begin by returning to Eq. (83). If we assume the slowly-varying amplitudes change with $y$ as $e^{iky}$ for some $k$, then the equations can be rearranged as:
@@
Given initial conditions $a_{L}(0)$ and $a_{R}(0)$ and our assumptions of codirectional and positive group velocities, Eq. (83) can be solved for the two amplitudes [63] to give:
@@
These equations are mathematically equivalent to those that govern optical couplers and switches [63, 130]. The acoustic power oscillates between the waveguides with a complete cycle (i.e. energy transferring back and forth) occurring over a wavelength of $\lambda_{\mathrm{trans}}=2\pi/k_{c}$. In terms of coupling rates this is equal to:
@@
If we assume that initially all the power is in one waveguide, say $a_{R}(0)=0$, then the maximum amount of power is in waveguide $R$ when $\cos(k_{c}y)=0$ and $\sin(k_{c}y)=\pm 1$. This occurs for example at $y=\pi/2k_{c}$, implying the maximum fraction of power that can be transferred is:
@@
The results of the simulation and our coupled modes expression are summarised in Table 1. The relative error between the analytic theory and simulation is generally $\lesssim 10\%$, and we can observe the qualitative behaviour predicted by the theory. Specifically, we see oscillation of power between the two waveguides. The two insets highlight the phase lag between the wavefronts predicted by Eq. (81). We can see for example in inset (i) that the bottom waveguide appears ‘ahead’ of the top waveguide at the position in space where there is a net power flow from it to the top waveguide. In contrast, inset (ii) shows how the relative phase between the waveguides is switched further down in the direction of travel, at a position where the net power flux is now from the top waveguide to the bottom waveguide.
@@ VIII Longitudinal evanescent coupling
A significant advantage of tunnel couplers is that they are directly compatible with highly-acoustically confining, highly-impedance mismatched suspended membranes. In contrast, coupling schemes like the parallel waveguides example in Section VII.1 rely on the lateral evanescent field and therefore must have a comparatively low impedance contrast. To achieve such a low impedance mismatch with suspended membranes would require a more complex fabrication process and quite possibly perform worse—for example, waveguides could be defined by depositing metal but this may introduce surface losses [24]. In other words, evanescent tunnel couplers are the natural solution for creating coupling in suspended membrane devices. They can also be used to create resonators with high (and tailorable) quality factors, as we will now discuss.
@@ IX Input-output formalism
Here $\gamma_{1}$ and $\gamma_{2}$ (with dimension $T^{-1}$) are the coupling rates to waveguides $1$ and $2$ respectively, and $\gamma_{0}$ is the intrinsic or internal loss rate describing coupling to the environment (which could for example correspond to air damping or thermoelastic damping in the membrane). The total loss rate $\gamma$ is equal to the sum of the couplings to the waveguides and the environment: $\gamma=\gamma_{0}+\gamma_{1}+\gamma_{2}$. The situation can be generalised to resonators coupled to $N$ waveguides by considering additional coupling rates $\gamma_{3},\,\gamma_{4},\,\ldots$ and so on. If the resonator is not being driven we can easily derive from Eq. (96) that the rate of energy loss is $\frac{\mathrm{d}|a|^{2}}{\mathrm{d}t}=-\gamma|a|^{2}$, consistent with our previous experience that $\gamma/2$ is the rate of amplitude decay.
@@
As illustrated in Fig. 28, the acoustic field in each waveguide is a superposition of waves travelling towards and away from the resonator. Because we assume the waveguides have infinite spatial extent, any waves travelling towards the resonator ($a_{\mathrm{{in,1}}},\,a_{\mathrm{{in,2}}}$) must originate from external sources; in contrast waves travelling away can either come from the resonator or be reflected inputs. It can be shown by applying time reversal symmetry that the loss rate for each waveguide also defines its reflection coefficient [140], such that the outgoing amplitude in each waveguide is simply: